### v1

# NuSTAR Detection of an Absorption Feature in ESP 39607: Evidence for an Ultra-fast Inflow?

Abstract
We report the serendipitous discovery of an absorption feature at 4.8 keV in the NuSTAR spectra of ESP 39607, a Seyfert 2 galaxy at , observed in May 2023 and August 2024. The feature is detected in both observations with individual significance levels between 2 and 3, computed with multiple statistical methods. The combined probability of detecting it in both observations is 4. The absorption feature is consistent with an ultra-fast inflow (UFI) potentially associated with Fe xxv or Fe xxvi K transitions. The inferred inflow velocity is 0.15–0.20, with an estimated launching radius of 22–89 , depending on the assumed iron transition and whether radiation pressure is accounted for. Photoionization modeling associates the UFI primarily with Fe xxv K absorption, blended with a minor contribution from Fe xxvi K. Alternative explanations, including associations with the warm-hot intergalactic medium or outflows of lighter elements, were investigated but found unlikely. If confirmed, this detection represents a rare example of a UFI, providing valuable evidence into extreme and/or non-standard accretion processes near supermassive black holes. Follow-up observations with higher-resolution X-ray spectroscopy, such as with XMM-Newton or XRISM, will be essential to confirm the nature of this feature and better constrain the physical mechanisms driving it.
1 Introduction
Active Galactic Nuclei (AGN) are the observational signature of supermassive black holes (SMBHs) accreting material at the centers of their host galaxies (e.g., Magorrian et al., 1998; Kormendy & Ho, 2013; Heckman & Best, 2014). Over the past two decades, X-ray observations have revealed peculiar absorption features in AGN spectra, which are commonly interpreted as signatures of winds or outflows driven by radiation pressure on ionized gas near the accretion disc (e.g., King & Pounds, 2015). In the most extreme cases, known as ultra-fast outflows (UFOs; Cappi et al., 2009; Tombesi et al., 2011, 2013; Matzeu et al., 2023, e.g.,), they reach velocities up to 0.1–0.5 (e.g., Chartas et al., 2002; Reeves et al., 2003; Braito et al., 2018; Serafinelli et al., 2019; Reeves et al., 2020; Braito et al., 2022; Gianolli et al., 2024). These powerful outflows play a critical role in AGN feedback processes, as the energy and momentum they carry can impact the surrounding interstellar medium in the host galaxy. This feedback mechanism is believed to regulate both SMBH growth and star formation, making UFOs a critical phenomenon in the co-evolution of SMBHs and their host galaxies (e.g., Cappi, 2006; King & Pounds, 2015; Laha et al., 2021; Lanzuisi et al., 2024).
Another class of phenomena that can be considered analogs of UFOs are ultra-fast inflows111We adopt the definition of “ultra-fast” for velocities from Tombesi et al., 2011. (UFIs), which remain considerably less studied despite their crucial role in understanding the inner accretion dynamics near SMBHs (e.g., King & Pringle, 2006; Pounds et al., 2018; Pounds & Page, 2024). For example, theoretical models and hydrodynamical simulations suggest that such inflows may originate from misaligned accretion discs that undergo tearing, potentially causing the disc to break and allowing material to fall inward rapidly under chaotic accretion (e.g., Nixon et al., 2012; Doǧan et al., 2018; Pounds et al., 2018). Thus, UFIs could offer valuable insights into these non-standard accretion processes, which typically assume cold accretion flows (e.g., Gaspari & Sadowski, 2017; Kobayashi et al., 2018), posing a challenge to traditional models of circular, planar accretion discs (e.g., Shakura & Sunyaev, 1973). Other possible origins for material infalling onto SMBHs include fallback from failed outflows, such as in the “aborted jet” or “failed wind” scenarios (e.g., Ghisellini et al., 2004; Proga & Kallman, 2004), where material launched from inner winds fails to escape and returns inward. UFIs, therefore, serve as valuable probes of these mechanisms and can help advance our understanding of accretion and feedback in the innermost regions of AGN.
Only a small number of documented cases of possible UFIs exist in the literature, including, e.g., Mrk 509 (Dadina et al., 2005), E1821+643 (Yaqoob & Serlemitsos, 2005), Mrk 335 (Longinotti et al., 2007), PG1211+143 (Reeves et al., 2005; Pounds et al., 2018; Pounds & Page, 2024), NGC 2617 (Giustini et al., 2017), and tentative detections from Tombesi et al. (2010). In addition, these inflow detections have typically been limited to low significance, with the most evident cases reaching no more than 3–3.5. This is because inflows not only share variability and unpredictability with outflows, but they also appear to be rarer (Laha et al., 2021). This rarity limits our understanding of the phenomenon, and key aspects, such as their identification and the physics behind what drives these events, remain unclear.
Here, we present the case study of ESP 39607, a highly luminous and heavily obscured Seyfert 2 galaxy at (Vettolani et al., 1998; Landi et al., 2010; Ricci et al., 2017), part of the BAT AGN Spectroscopic Survey (BASS DR2; Oh et al., 2018; Koss et al., 2022a, b). In this source, we report the detection of an absorption feature in two NuSTAR observations, which we tentatively identify as a UFI.
The paper is organized as follows: in Section §2, we describe the data used in this work; in Section §3, we present our analysis; and in Section §4, we discuss the main results and their implications. Throughout this work, we assumed a CDM cosmology with the fiducial parameters km s-1 Mpc-1, , and . Uncertainties are reported at a 90% confidence level, if not stated otherwise.
2 Data
ESP 39607 was observed with NuSTAR for 21.2 ks in May 2023, and for 21.6 ks in August 2024 as a follow-up of the first observation. The data were reduced using NuSTARDAS v2.1.2. Specifically, we calibrated and cleaned the downloaded datasets using the standard nupipeline routine. The spectral extraction was done using the nuproducts command for the two FPMA and FPMB cameras. We extracted the source spectra from a circular region with a radius of 60″. The background was extracted from two 75″ circular regions close to the source region, on the same detector, avoiding chip gaps, and ensuring no other sources were nearby. All the spectra were binned with a minimum of 5 counts/bin, balancing information loss and possible biases in W-statistic fitting (Buchner & Boorman, 2023).
For the first observation, no background flares or other issues that could significantly impact the data were identified. For the second observation, the NuSTAR background report222https://nustarsoc.caltech.edu/NuSTAR_Public/
NuSTAROperationSite/SAA_Filtering/saa_reports/ flagged the source for solar flare activity. However, the impact on the data was minimal (NuSTAR, private communication). We adopted a conservative approach by following the NuSTAR recommended guidelines for identifying and mitigating solar flare contamination333https://github.com/NuSTAR/nustar-gen-utils. Specifically, we used the GTI_filter_solar_flare.ipynb notebook to create custom good time intervals (GTIs) for use in the nuproducts routine. This resulted in the removal of less than 1% of the total exposure from the initial extraction, confirming the minimal impact of flaring activity on the observation.
In addition to the NuSTAR data analyzed in this work, a 60.7 ks Suzaku observation from December 2010 is also available for ESP 39607. We used the spectra extracted by Ricci et al. (2017) and refer the reader to that work for details on data reduction. The same dataset is also included and shown in the broad-band spectral fitting analysis presented in Peca et al. (submitted), and we refer to both studies for additional modeling details. Applying the same procedure described in Section 3, we did not find evidence of the absorption feature seen in the NuSTAR spectra within the Suzaku data. The implications of this observation are discussed in Section 4.
3 Spectral Analysis
The spectral analysis was performed with XSPEC v12.14.0 (Arnaud, 1996). W-statistic (Cash, 1979; Wachter et al., 1979) was applied. All the fits were performed in the 3–35 keV energy range, as at higher energies the background dominates over the source emission.
3.1 May 2023 observation
We started by fitting the 2023 observation, which has a net (i.e., background subtracted) count rate of 0.0500.002 . First, we modeled the source spectrum using a simple model composed of an absorbed power-law with an energy cut-off fixed at 200 keV (e.g., Ricci et al., 2017; Zappacosta et al., 2018) plus a pexrav (Magdziarz & Zdziarski, 1995) component to account for reflected emission. For simplicity, we assumed pure reflection by fixing (e.g., Piconcelli et al., 2011; Zhao et al., 2019; Serafinelli et al., 2024); the negative value directs the model to operate in reflection-only mode (e.g., Alexander et al., 2011; Boorman et al., 2025), while a value of one corresponds to reflection from an infinite slab illuminated by isotropic coronal emission (e.g., Marchesi et al., 2016; Zappacosta et al., 2018). The inclination angle was set to the default value of 60∘ (e.g., Ueda et al., 2007; Zhao et al., 2020). After verifying consistency between the normalizations of the primary power-law and reflection components, we linked them to the same value to minimize degeneracies (e.g., Liu et al., 2017; Peca et al., 2023b). We refer to this configuration as Model 1. Second, we fitted the source spectrum with the more complex, but physically motivated and self-consistent model, UXCLUMPY (Buchner et al., 2019). We fixed the TORsigma and CTKcover parameters to their standard values, 28 and 0.4, respectively, and the inclination angle to 60∘, as for the first model. We refer to this configuration as Model 2. Results from Model 2 are consistent with those from Model 1 within the uncertainties (see Table 1), and both models confirm the obscured and luminous nature of the source, with derived cm-2 and erg s-1, in agreement with those previously reported by Ricci et al. (2017).
The fit residuals revealed a possible absorption feature at 4.8 keV (Figure 1, left panel). To model this feature, we introduced a Gaussian line component to both models, where line normalization and energy were left free to vary. The line width was fixed to eV (e.g., Tombesi et al., 2010; Vignali et al., 2015), as it was found to be unconstrained at the lower end when left free, i.e., consistent with a narrow and unresolved feature. This inclusion improved the fit significantly. For Model 1, we obtained /2, which corresponds roughly to a fit improvement of 3 assuming Gaussian approximation (e.g., Tozzi et al., 2006). For Model 2, the improvement was /2, i.e., 2.7. Model 1 yielded keV and an equivalent width of keV, with Model 2 giving consistent results. We summarize all the detailed results for all model configurations in Table 1.
We performed an Akaike Information Criterion test (AIC, Akaike, 1974) to compare models with and without the Gaussian line. In brief, AIC provides a relative estimate of model quality by balancing goodness of fit against model complexity, and it is valid for both nested and non-nested models. It is defined as , where is the maximum likelihood from the fit and is the number of free parameters. The model with the lowest AIC is preferred. According to the Burnham & Anderson (2002) scale, indicates that the model without the additional Gaussian line has “considerably less” statistical support, while suggests the support for the model without the line is “essentially none”. Here, we interpret as providing “positive” evidence, as “strong” evidence, and as “very strong” evidence in favor of the model with the Gaussian line. We obtained for Model 1 and for Model 2, thereby statistically justifying the inclusion of the Gaussian line in the preferred model as strong and positive, respectively.
To further assess the reliability of the absorption feature detection, we adopted a Monte Carlo (MC) approach, as widely used in the literature to evaluate the significance of emission and absorption lines (e.g., Markowitz et al., 2006; Vignali et al., 2015; Costanzo et al., 2022; Peca et al., 2023a). We simulated spectra using the best-fit models without the Gaussian line, keeping the same responses and background of the observed data. We then fitted each simulated spectrum by adding a Gaussian line and performing a scan over the 3–10 keV energy range using the XSPEC steppar command with steps of keV. This range was chosen to avoid the drop in NuSTAR ’s effective area at 10 keV, which could compromise the reliability of the process (Peca et al., 2021). The resulting best-fit statistic is compared to the statistic from the model without the absorption line. This allows us to calculate the fraction of cases where the exceeds that found in the real spectrum, i.e., the fraction of cases where a fluctuation in the spectrum has a significance equal or higher than that of the absorption line in the real spectrum. This occurred in only 1.4% of the simulations for Model 1 and 5.0% for Model 2. In other words, the probability that the absorption feature is real is 98.6% for Model 1 and 95.0% for Model 2, which can be interpreted as detection significances of approximately 2.5 and 2.0, respectively.
3.2 August 2024 observation
Based on these results, a follow-up NuSTAR observation was conducted in August 2024. The August 2024 observation has a much lower net count rate, 0.0330.002 . Compared to the 2023 observation, this indicates a decrease in flux of 34% and in signal-to-noise ratio444Computed as , where and are the source and background counts in the source region. from 30 to 24. We performed the same spectral analysis as described for the 2023 observation. The spectrum and confidence contours are shown in Figure 2. The new observation confirmed the absorption feature, as adding a Gaussian line improved the fit by /2 and 11.1/2 for Model 1 and Model 2, respectively, corresponding to significance levels of and . The AIC criterion further supports the presence of the absorption line, yielding and , indicating a “strong” statistical evidence in favor of its inclusion in both models.
Interestingly, compared to the 2023 observations, the best-fit line energy is shifted to a slightly lower energy of keV. However, we note that the line centroids are consistent within the 99% confidence level, and the observed difference is also compatible with NuSTAR ’s spectral resolution at these energies (400 eV). Compared to the earlier observation, the absorption feature also shows a slightly higher equivalent width, though consistent within the uncertainties. No significant variation in other model parameters was found between the two NuSTAR observations.
We further explored the scenario where the absorption feature is the same between the two observations by performing MC-based simulations. Starting from the best-fit models of the 2023 observation, including the absorption line, we simulated spectra after adjusting the model normalization to match the lower flux and signal-to-noise ratio of the 2024 observation. Each simulated spectrum was then fitted using the same procedure described earlier, using the steppar command to scan the spectra and identify the line centroid. The resulting distribution is shown in Figure 3. The 4.5 keV centroid measured in the 2024 observation falls well within the 5th percentile of the simulated distributions for both models (4.2 keV for Model 1 and 4.0 keV for Model 2), and is very close to the 16th percentile (4.5 keV for both models), corresponding to the 90% and 68% confidence levels, respectively. We therefore conclude that, from a purely statistical standpoint, the absorption line in the 2024 observation is consistent with being the same as that in 2023.
3.3 Simultaneous fitting
We then performed a simultaneous fit of the 2023 and 2024 observations. To account for the decrease in flux in the 2024 dataset, we introduced a multiplicative constant to allow for a different normalization between the two observations. For both models, this constant was found to be when including the Gaussian line, and when it was not. Since the other model parameters were found to be consistent when fitting the observations independently, we linked them across the two datasets. For the absorption line, however, we allowed the energy and normalization to vary freely between the two observations. This choice was motivated by the fact that linking the line component forces the fit to settle between the two values observed in the individual fits, failing to properly capture the line profiles in either epoch. Indeed, when the line energy was linked, we obtained a centroid of keV, with values of 13.4/2 and 10.7/2 for Model 1 and Model 2, respectively, which are fit improvements comparable to those obtained when fitting the observations separately. Instead, when allowing the line energy and normalization to vary independently, the fit significantly improved. We found values of 25.8/4 and 21.9/4 for Model 1 and Model 2, respectively, corresponding to significances of 4.1 and 3.7. The associated AIC values are 17.8 and 13.9, indicating a “very strong” preference for the models with the Gaussian line. The best-fit values for the other model parameters remain consistent with those obtained from the individual fits and are reported in Table 1.
To further quantify the significance of detecting the same absorption feature in both observations, we performed MC-based simulations following the same procedure used for the individual epochs. Starting from the best-fit continuum models (i.e., without any absorption line), we simulated spectra for each observation. For each simulated pair, we searched for spurious absorption features in each spectrum with a equal to or greater than that observed when including the Gaussian line. We then required that these features appear in both spectra within an energy separation of 0.3 keV, corresponding to the difference between the observed centroids. The resulting false positive probabilities translate into detection significances of 4.4 for Model 1 and 4.2 for Model 2, reinforcing the robustness of the detection across both epochs.
Obs. | Model | EW | cstat/dof | cstat | AIC | AIC | MC | ||||
2023 | 1 | 34.0 | 1.95 | 4.2 | - | - | 368.8/378 | 374.8 | |||
2023 | 1+Line | 34.1 | 1.98 | 4.6 | 4.8 | -0.22 | 356.6/376 | 3.0 | 366.6 | strong | 2.5 |
2023 | 2 | 26.8 | 1.82 | 6.2 | - | - | 367.9/378 | 373.9 | |||
2023 | 2+Line | 27.2 | 1.90 | 6.7 | 4.8 | -0.24 | 357.8/376 | 2.7 | 367.8 | positive | 2.0 |
2024 | 1 | 37.2 | 1.94 | 2.8 | - | - | 274.7/279 | 280.7 | |||
2024 | 1+Line | 29.7 | 1.90 | 2.6 | 4.5 | -0.31 | 261.6/277 | 3.2 | 271.6 | strong | 2.4 |
2024 | 2 | 31.2 | 1.81 | 4.5 | - | - | 272.9/279 | 278.9 | |||
2024 | 2+Line | 24.7 | 1.77 | 3.9 | 4.5 | -0.27 | 261.8/277 | 2.9 | 271.8 | strong | 2.2 |
Both | 1 | 34.7 | 1.96 | 4.2∗ | - | - | 645.0/659 | 653.0 | |||
Both | 1+Line | 32.3 | 1.99 | 4.3∗ | 4.8 | -0.20 | 619.2/655 | 4.1 | 635.2 | very strong | 4.4 |
4.5 | -0.32 | ||||||||||
Both | 2 | 27.4 | 1.80 | 6.1∗ | - | - | 642.5/659 | 650.5 | |||
Both | 2+Line | 25.9 | 1.83 | 6.2∗ | 4.8 | -0.20 | 620.6/655 | 3.7 | 636.6 | very strong | 4.2 |
4.5 | -0.33 |
4 Discussion
4.1 A possible ultra-fast inflow?
We first explore the interpretation that the observed absorption feature arises from an ionized inflow, a plausible physical scenario for an absorption line at these energies (e.g., Longinotti et al., 2007; Tombesi et al., 2010; Pounds et al., 2018; Pounds & Page, 2024). At the source redshift (), the observed 4.8 keV feature corresponds to 5.8 keV in the rest frame. Given the observed energy range, the most likely candidates are Fe xxv (rest-frame, 6.70 keV) or Fe xxvi (6.97 keV) K transitions, with iron favored due to its high cosmic abundance compared to other elements and frequent detection in X-ray spectra (e.g., Reeves et al., 2005; Longinotti et al., 2007). We focus exclusively on the K transition, as a reliable identification of K would require simultaneous detection of both transitions, which we do not observe. For the following analysis, we rely on the 2023 observation, which provides the highest signal-to-noise ratio.
4.1.1 Inflow velocity and position
We estimated the inflow velocity using the the relation (e.g., Tombesi et al., 2011; Serafinelli et al., 2023; Bertola et al., 2020), where is the observed redshift of the absorber, is the intrinsic absorber redshift, and is the cosmological redshift of the source. The inflow velocity is then derived from , where . This calculation indicates an ultra-fast inflow (UFI) with a velocity of 0.20 when assuming Fe xxvi K and 0.15 assuming Fe xxv K, consistent with previous UFI studies (e.g., Dadina et al., 2005; Giustini et al., 2017; Pounds et al., 2018). We note that, given the limited spectral resolution of our data (400 eV), we cannot distinguish between these two lines, and therefore the detected feature could be a blend of both transitions.
To test the physical consistency of these findings, we compared the inferred inflow velocity with a simple model describing radially infalling gas under gravitational attraction, influenced by radiation pressure, following Longinotti et al. (2007). In this scenario, we calculate the minimum radius at which the observed inflow velocity equals the gravitational free-fall velocity. This radius can be expressed as
where is the gravitational constant, is the inflow velocity, is the black hole mass, and is the radiation pressure. is derived from the average X-ray luminosity between the two used models, using the bolometric correction of Duras et al. (2020). For Fe xxvi K (Fe xxv K), this yields , where is the gravitational radius. When radiation pressure is neglected, the corresponding values are . These distances are consistent with those found in the literature for UFIs (e.g., Longinotti et al., 2007; Pounds et al., 2018). Assuming a black hole mass of , derived from the velocity dispersion method described in Koss et al. (2022c) and included in the forthcoming BASS DR3 (Koss et al., in prep.), the corresponding physical distances are pc and pc for the two cases, respectively.
4.1.2 Photoionization modelling
We then employed a more physically motivated approach by modeling the absorption feature using the XSTAR v2.59 photoionization code (Bautista & Kallman, 2001; Kallman & Bautista, 2001). Following Tombesi et al. (2011), we assumed a power-law with , gas temperature T=106K (Nicastro et al., 1999; Bianchi et al., 2005), and a gas density of cm-3. Solar abundances were adopted from Asplund et al. (2009). We simulated three grids555We used the MPI C++ parallelization tool provided by Danehkar et al. (2018) to run XSTAR. with different turbulence velocities, 100, 1000 and 5000 , and a fixed covering factor of 1 (e.g., Pounds et al., 2018). The additional parameters given by these grids are the column density of the absorbing material (), the redshift of the absorber (), and the ionization parameter (). The latter is defined by (erg s-1 cm), where is the ionizing luminosity of the source integrated between 1 and 1000 Ryd (1 Ryd = 13.6 eV), is the gas density, and is the distance between the absorber and the central source (Tarter et al., 1969).
We obtained the best results using the XSTAR grid with a turbulence velocity of 5000 , which provided both constrained parameters and the most improvement to the fit. The results are summarized in Table 2, and our best-fit result for Model 1 is shown in Figure 4. For both models we obtained consistent results, with an ionization parameter and an absorbing column density for the inflowing material of cm-2, with a redshift of , corresponding to an inflow velocity of 0.15. According to the XSTAR modeling, the absorption feature at this redshift is primarily produced by the Fe xxv K transition. As shown in Figure 4, the observed absorption feature is reproduced by the model as a blend of Fe xxv and Fe xxvi K lines, with Fe xxv K providing the dominant contribution. The inclusion of the XSTAR component improves the fit by cstat/d.o.f. = 7.8/3 and 8.3/3 for Models 1 and 2, respectively, corresponding to confidence for three additional free parameters.
For comparison, the 1000 grid produced results broadly consistent with the 5000 case, but with larger uncertainties and poorer fits (cstat/d.o.f. = 5.0/3 and 6.3/3 for Models 1 and 2, respectively), as it failed to model the full depth of the absorption feature. For the same reason, the 100 grid did not allow us to constrain any of the absorber parameters.
Obs. | Model | cstat/dof | cstat | ||||||
2023 | 1+XSTAR | 35.7 | 1.99 | 5.0 | 19 | 3.0 | 0.39 | 361.0/375 | 2.0 |
2023 | 2+XSTAR | 27.6 | 1.96 | 7.5 | 19 | 3.1 | 0.39 | 359.6/375 | 2.1 |
4.1.3 On the Suzaku 2010 observation
As described in Section 2, an archival Suzaku observation from 2010 is available for ESP 39607. These data were initially reduced by Ricci et al. (2017) as part of a broad AGN characterization study, and have been included in a multi-epoch AGN analysis by Peca et al. (submitted), which also incorporates the NuSTAR observations presented here. In particular, the Peca et al. analysis confirms the spectral results reported in this work and classifies the source as variable in flux, with the 2010 Suzaku observation showing a 2–10 keV flux 20% higher than that of the 2023 NuSTAR observation. Peca et al. also show that soft X-ray components, such as thermal emission and scattering from the primary power-law emission, are too faint to affect the hard X-ray spectra analyzed here, with negligible impact above 3 keV in the observed frame. We refer the reader to this work for further details.
Applying the same analysis and fitting procedure outlined in Section 3, we found no significant evidence of the absorption feature in the Suzaku data. The most plausible explanation for this absence is the intrinsic variability of the infalling material. Indeed, observational studies have shown that inflows and outflows commonly vary on timescales ranging from days to years (e.g., Tombesi et al., 2010; Pounds et al., 2018; Laha et al., 2021; Braito et al., 2022), which is not at odds with the lack of detection in the 2010 observation and supports the interpretation of the feature as a transient episode of UFI.
4.1.4 The physical mechanism that drives the inflow
Several physical mechanisms could explain the origin of the transient UFI detected in this source. One possibility is that the inflowing material results from instabilities in a misaligned accretion disk. Theoretical models and hydrodynamical simulations suggest that such disks can undergo tearing and fragmentation, causing portions of the disk to break apart into rings or clumps that decouple from the main flow (e.g., Nixon et al., 2012; Doǧan et al., 2018). These fragments can then fall inward rapidly on plunging or eccentric trajectories, leading to short-lived inflow events. This form of chaotic accretion deviates from the standard picture of smooth, circular disk accretion (e.g., Shakura & Sunyaev, 1973), and is typically associated with cold, transient events (e.g., Gaspari & Sadowski, 2017; Kobayashi et al., 2018).
Alternatively, the inflow could result from material that was initially launched outward but failed to escape the gravitational pull of the central SMBH (e.g., Giustini et al., 2017). In the “aborted jet” scenario (Ghisellini et al., 2004), compact blobs of plasma are intermittently ejected from the inner accretion flow but lack the energy to fully escape. These blobs may rise to a certain distance before falling back, potentially interacting with new ejecta and producing shocks contributing to the X-ray emission. A related possibility is the “failed wind” scenario (Proga & Kallman, 2004), in which the absorbing material is part of an ongoing wind that becomes over-ionized by the intense continuum radiation. As the wind loses momentum, it stalls and eventually falls back toward the disk plane. In both cases, the inflow is expected to be highly variable and short-lived, with characteristic timescales on the order of the dynamical or free-fall time.
Under these scenarios, the emergence of the absorption feature in the 2023/2024 NuSTAR data, following its absence in the 2010 Suzaku observation, is consistent with the expected episodic nature of such events, which are predicted to vary on timescales of years or less. Additionally, the accompanying flux variations of 20% between the 2010 Suzaku and 2023 NuSTAR observations, and 34% between the 2023 and 2024 NuSTAR observations, further support these interpretations, as such changes are likely to be expected for disk instabilities or fallback processes (e.g., Ghisellini et al., 2004; Nixon et al., 2012).
4.2 Other scenarios
While our primary interpretation of the observed feature is an ionized iron absorption line from inflowing material, a scenario supported by consistency with previous UFI studies (e.g., Dadina et al., 2005; Reeves et al., 2005; Giustini et al., 2017; Pounds et al., 2018), in this Section we consider alternative potential explanations.
One possibility is that the absorption originates from the Warm-Hot Intergalactic Medium (WHIM). However, its variability between the two NuSTAR observations and absence in the Suzaku data is inconsistent with the typically stable nature of WHIM absorption (Nicastro et al., 2018, 2023). Moreover, the column densities required to reproduce the observed feature ( cm-2) are much larger than those typically associated with diffuse local gas ( cm-2; Zappacosta et al., 2010; Nicastro et al., 2018, 2023, e.g.,). An absorption line in the 4–5 keV range in the spectrum of an AGN at is also unlikely to originate from the WHIM. Indeed, the most abundant WHIM tracers in absorption, such as O vii, O viii, and Ne ix (e.g., Kaastra et al., 2006; Nicastro et al., 2018, 2023), produce absorption features at softer X-ray energies ( keV). These considerations make it unlikely that the observed feature is related to the WHIM.
Another potential explanation is that the feature arises from an outflow involving elements lighter than iron (e.g., Gupta et al., 2014; Reeves et al., 2020; Lanzuisi et al., 2024). While possible in principle, most lighter elements produce K transitions at significantly lower energies, such as Ca xx (rest-frame, 4.11 keV), Si xiv (2.00 keV), S xvi (2.62 keV), and Mg xii (1.47 keV). To explain an absorption line observed at 4.8 keV, the required outflow velocities for S xvi, and Mg xii ions would need to exceed 0.75, making this scenario highly unlikely. Si xiv would require an outflow velocity of 0.65, a velocity that has only been observed in a handful of exceptional sources and exclusively associated with highly ionized iron transitions (e.g., Chartas et al., 2009; Lanzuisi et al., 2012; Luminari et al., 2023). The only ion among these with a rest-frame energy close enough to match the observed line with a more moderate outflow velocity is Ca xx, which would require a velocity of 0.33. However, calcium is 15 times less abundant than iron (Asplund et al., 2009), making a strong, isolated Ca xx absorption feature difficult to justify. To further test the possibility of outflowing elements lighter than iron, we performed a blind velocity search using the XSTAR model, allowing the redshift parameter to vary into the negative velocity regime (i.e., outflows). We tested both the 5000 and 1000 turbulence grids, using the steppar command with a step of . No significant improvement in fit statistics was found, thus providing no support for this scenario.
Based on these analyses, we conclude that the UFI interpretation is the most plausible explanation for the observed absorption feature, although we cannot entirely rule out alternative scenarios with the current data.
5 Summary and conclusions
We presented the detection and analysis of an absorption feature in the X-ray spectra of ESP 39607, a Seyfert 2 galaxy at , observed with NuSTAR in May 2023 and August 2024. Based on our best-fit modeling, we associate this feature with an ultra-fast inflow (UFI) at a velocity of 0.15, primarily produced by the Fe xxv K transition, with a blended contribution from Fe xxvi K. The observed 4.8 keV feature corresponds to 5.8 keV in the rest frame of the source. Below, we summarize our findings:
The absorption feature was consistently detected at 4.8 keV in both 2023 and 2024 NuSTAR observations. Although the second observation yielded a slightly different best-fit energy (4.5 keV), simulations demonstrate this shift is not statistically significant, reinforcing the consistency of the detections. Individually, each detection has a significance that ranges between 2 and 3, as estimated through MC simulations and other statistical tests. When the two observations are considered together, the significance exceeds 4.
The feature was first fitted with a Gaussian absorption line, combined with the two different continuum models that describe the rest of the spectrum. Based on the best fits, the line is tentatively interpreted as a redshifted Fe xxv K or Fe xxvi K transition, corresponding to an inflow velocity of 0.15–0.20.
A simple spherical model for radially infalling gas suggests that the inflow originates at 22 (34) for Fe xxvi (Fe xxv), or at 55 (89) if radiation pressure is considered. These estimates are consistent with values reported in the literature.
When fitting the data with XSTAR photoionization models, the best-fit was obtained using a grid with a turbulence velocity of 5000 , where the absorption feature is dominated by a strong Fe xxv K line, blended with a minor contribution from Fe xxvi. The derived parameters, , cm-2, and an inflow velocity of 0.15, are overall in agreement with previous UFIs studies.
We tested alternative interpretations to explain the origin of the absorption feature, including the presence of WHIM signatures and outflows from lighter elements. Based on these analyses, we conclude that the ultra-fast inflow interpretation is the most plausible explanation.
Although our results favor an inflow interpretation, we acknowledge the limitations of the current data. While NuSTAR enabled the detection of the absorption feature, its 400 eV energy resolution prevents distinguishing between closely spaced lines such as Fe xxv and Fe xxvi, and limits our ability to characterize the absorber in greater detail. To address this, we have secured observing time with XMM-Newton (AO 24, PI A. Peca), which will provide significantly improved sensitivity and spectral resolution. Additionally, the new XRISM (XRISM Science Team, 2020) mission will be instrumental in further probing the nature and variability of this feature, and more broadly, in understanding the role of UFIs in AGN accretion and feedback.

### v2

@@
 Photoionization modeling associates the UFI primarily with Fe xxv K absorption, blended with a minor contribution from Fe xxvi K.
-Alternative explanations, including associations with the warm-hot intergalactic medium or outflows of lighter elements, were investigated but found unlikely.
+Alternative explanations, including associations with the warm-hot intergalactic medium or outflows of lighter elements, are investigated but found unlikely.
 If confirmed, this detection represents a rare example of a UFI, providing valuable evidence into extreme and/or non-standard accretion processes near supermassive black holes.
@@
 This resulted in the removal of less than 1% of the total exposure from the initial extraction, confirming the minimal impact of flaring activity on the observation.
-In addition to the NuSTAR data analyzed in this work, a 60.7 ks Suzaku observation from December 2010 is also available for ESP 39607.
+In addition to the NuSTAR data analyzed in this work, a 60.7 ks Suzaku/XIS observation from December 2010 is also available for ESP 39607.
 We used the spectra extracted by Ricci et al.
@@
 The implications of this observation are discussed in Section 4.
+A summary of the observations analyzed in this work is provided in Table 1.
+Telescope | Obs. ID | Date | Exp. | Flux |
+NuSTAR | 60860001002 | 2023-05-08 | 21.2 | 1.55 |
+NuSTAR | 60161025002 | 2024-08-22 | 21.6 | 1.01 |
+Suzaku | 705048010 | 2010-12-19 | 60.7 | 1.89 |
 3 Spectral Analysis
@@
 For simplicity, we assumed pure reflection by fixing (e.g., Piconcelli et al., 2011; Zhao et al., 2019; Serafinelli et al., 2024); the negative value directs the model to operate in reflection-only mode (e.g., Alexander et al., 2011; Boorman et al., 2025), while a value of one corresponds to reflection from an infinite slab illuminated by isotropic coronal emission (e.g., Marchesi et al., 2016; Zappacosta et al., 2018).
-The inclination angle was set to the default value of 60∘ (e.g., Ueda et al., 2007; Zhao et al., 2020).
+The inclination angle was set to the default value of 60∘ (e.g., Ueda et al., 2007; Zhao et al., 2020), and the photon index was linked to that of the primary power-law component.
 After verifying consistency between the normalizations of the primary power-law and reflection components, we linked them to the same value to minimize degeneracies (e.g., Liu et al., 2017; Peca et al., 2023b).
@@
 The line width was fixed to eV (e.g., Tombesi et al., 2010; Vignali et al., 2015), as it was found to be unconstrained at the lower end when left free, i.e., consistent with a narrow and unresolved feature.
-This inclusion improved the fit significantly.
+We used the XSPEC steppar command to map the confidence contours on line normalization and energy, and show them in Figure 2.
+The inclusion of this absorption line improved the fit significantly.
 For Model 1, we obtained /2, which corresponds roughly to a fit improvement of 3 assuming Gaussian approximation (e.g., Tozzi et al., 2006).
@@
 We simulated spectra using the best-fit models without the Gaussian line, keeping the same responses and background of the observed data.
-We then fitted each simulated spectrum by adding a Gaussian line and performing a scan over the 3–10 keV energy range using the XSPEC steppar command with steps of keV.
+We then fitted each simulated spectrum by adding a Gaussian line and performing a scan over the 3–10 keV energy range using the steppar command with steps of keV.
 This range was chosen to avoid the drop in NuSTAR ’s effective area at 10 keV, which could compromise the reliability of the process (Peca et al., 2021).
@@
 The August 2024 observation has a much lower net count rate, 0.0330.002 .
-Compared to the 2023 observation, this indicates a decrease in flux of 34% and in signal-to-noise ratio444Computed as , where and are the source and background counts in the source region.
+Compared to the 2023 observation, this indicates a decrease in flux of 35% and in signal-to-noise ratio444Computed as , where and are the source and background counts in the source region.
 from 30 to 24.
 We performed the same spectral analysis as described for the 2023 observation.
-The spectrum and confidence contours are shown in Figure 2.
+The spectrum and confidence contours are shown in Figure 1 (right panel) and Figure 2, respectively.
 The new observation confirmed the absorption feature, as adding a Gaussian line improved the fit by /2 and 11.1/2 for Model 1 and Model 2, respectively, corresponding to significance levels of and .
@@
 The best-fit values for the other model parameters remain consistent with those obtained from the individual fits and are reported in Table 2.
-To further quantify the significance of detecting the same absorption feature in both observations, we performed MC-based simulations following the same procedure used for the individual epochs.
-Starting from the best-fit continuum models (i.e., without any absorption line), we simulated spectra for each observation.
-For each simulated pair, we searched for spurious absorption features in each spectrum with a equal to or greater than that observed when including the Gaussian line.
-We then required that these features appear in both spectra within an energy separation of 0.3 keV, corresponding to the difference between the observed centroids.
-The resulting false positive probabilities translate into detection significances of 4.4 for Model 1 and 4.2 for Model 2, reinforcing the robustness of the detection across both epochs.
 Obs. | Model | EW | cstat/dof | cstat | AIC | AIC | MC | ||||
@@
 4.5 | -0.33 |
+To further quantify the significance of detecting the same absorption feature in both observations, we performed MC-based simulations following the same procedure used for the individual epochs.
+Starting from the best-fit continuum models (i.e., without any absorption line), we simulated spectra for each observation.
+For each simulated pair, we searched for spurious absorption features in each spectrum with a equal to or greater than that observed when including the Gaussian line.
+We then required that these features appear in both spectra within an energy separation of 0.3 keV, corresponding to the difference between the observed centroids.
+The resulting false positive probabilities translate into detection significances of 4.4 for Model 1 and 4.2 for Model 2, reinforcing the robustness of the detection across both epochs.
 4 Discussion
@@
 with different turbulence velocities, 100, 1000 and 5000 , and a fixed covering factor of 1 (e.g., Pounds et al., 2018).
-The additional parameters given by these grids are the column density of the absorbing material (), the redshift of the absorber (), and the ionization parameter ().
+The additional free parameters given by these grids are the column density of the absorbing material (), the redshift of the absorber (), and the ionization parameter ().
 The latter is defined by (erg s-1 cm), where is the ionizing luminosity of the source integrated between 1 and 1000 Ryd (1 Ryd = 13.6 eV), is the gas density, and is the distance between the absorber and the central source (Tarter et al., 1969).
@@
 = 7.8/3 and 8.3/3 for Models 1 and 2, respectively, corresponding to confidence for three additional free parameters.
+The slightly lower significance, compared to the models with the Gaussian line in absorption, arises because the XSTAR grid introduces three free parameters (, , and ) versus only two for the Gaussian line (line energy and normalization).
+Additionally, XSTAR accounts for all resonance absorption lines and edges from elements up to in multiple ionization states, among which only the feature presented in this work is statistically significant in our NuSTAR dataset.
 For comparison, the 1000 grid produced results broadly consistent with the 5000 case, but with larger uncertainties and poorer fits (cstat/d.o.f.
@@
 Under these scenarios, the emergence of the absorption feature in the 2023/2024 NuSTAR data, following its absence in the 2010 Suzaku observation, is consistent with the expected episodic nature of such events, which are predicted to vary on timescales of years or less.
-Additionally, the accompanying flux variations of 20% between the 2010 Suzaku and 2023 NuSTAR observations, and 34% between the 2023 and 2024 NuSTAR observations, further support these interpretations, as such changes are likely to be expected for disk instabilities or fallback processes (e.g., Ghisellini et al., 2004; Nixon et al., 2012).
+Additionally, the accompanying flux variations of 20% between the 2010 Suzaku and 2023 NuSTAR observations, and 35% between the 2023 and 2024 NuSTAR observations, further support these interpretations, as such changes are likely to be expected for disk instabilities or fallback processes (e.g., Ghisellini et al., 2004; Nixon et al., 2012).
 4.2 Other scenarios
@@
 The absorption feature was consistently detected at 4.8 keV in both 2023 and 2024 NuSTAR observations.
-Although the second observation yielded a slightly different best-fit energy (4.5 keV), simulations demonstrate this shift is not statistically significant, reinforcing the consistency of the detections.
+Although the second observation yielded a slightly different best-fit energy (4.5 keV), simulations demonstrate that this shift is not statistically significant, reinforcing the consistency of the detections.
 Individually, each detection has a significance that ranges between 2 and 3, as estimated through MC simulations and other statistical tests.
 When the two observations are considered together, the significance exceeds 4.
-The feature was first fitted with a Gaussian absorption line, combined with the two different continuum models that describe the rest of the spectrum.
+The absorption feature was first fitted with a Gaussian absorption line, combined with two different continuum models that describe the rest of the spectrum.
+Both models yielded consistent results.
 Based on the best fits, the line is tentatively interpreted as a redshifted Fe xxv K or Fe xxvi K transition, corresponding to an inflow velocity of 0.15–0.20.
@@
 These estimates are consistent with values reported in the literature.
-When fitting the data with XSTAR photoionization models, the best-fit was obtained using a grid with a turbulence velocity of 5000 , where the absorption feature is dominated by a strong Fe xxv K line, blended with a minor contribution from Fe xxvi.
+When fitting the data with XSTAR photoionization models, the best-fit was obtained using a grid with a turbulence velocity of 5000 , where the absorption feature is dominated by a strong Fe xxv K line, blended with a minor contribution from Fe xxviK.
 The derived parameters, , cm-2, and an inflow velocity of 0.15, are overall in agreement with previous UFIs studies.
@@
 Although our results favor an inflow interpretation, we acknowledge the limitations of the current data.
-While NuSTAR enabled the detection of the absorption feature, its 400 eV energy resolution prevents distinguishing between closely spaced lines such as Fe xxv and Fe xxvi, and limits our ability to characterize the absorber in greater detail.
-To address this, we have secured observing time with XMM-Newton (AO 24, PI A.
-Peca), which will provide significantly improved sensitivity and spectral resolution.
-Additionally, the new XRISM (XRISM Science Team, 2020) mission will be instrumental in further probing the nature and variability of this feature, and more broadly, in understanding the role of UFIs in AGN accretion and feedback.
+While NuSTAR enabled the detection of the absorption feature, its 400 eV energy resolution prevents distinguishing between closely spaced lines such as Fe xxv and Fe xxvi K transitions, and limits our ability to characterize the absorber in greater detail.
+To address this limitation, we have secured observing time with XMM-Newton (AO 24, PI A.
+Peca).
+This XMM-Newton observation is expected to provide significantly improved sensitivity and spectral resolution ( eV), making it ideally suited for confirming the absorption feature and better constraining the blended Fe xxv and Fe xxvi.
+Furthermore, the newly launched XRISM mission (XRISM Science Team, 2020), with its unprecedented energy resolution (5 eV), could enable detailed characterization of the absorber.
+This would be especially valuable if the intrinsic turbulent velocity is lower than measured here, potentially allowing XRISM to resolve the individual Fe xxv and Fe xxvi components.
+Together, these missions will offer complementary insights into the nature and variability of this feature, thereby improving our broader understanding of the role that UFIs play in AGN accretion and feedback processes.
+acknowledgments
+We acknowledge the anonymous referee for the valuable comments that improved the quality of the paper.
+A.P.
+acknowledges D.
+Costanzo for the useful discussions.
+A.P.
+and M.K.
+thank the NuSTAR team for the helpful discussions on the observations and the data reduction of ESP 39607.
+A.P.
+and M.K.
+acknowledge support from NASA through ADAP award 80NSSC22K1126 and NuSTAR grants 80NSSC22K1933 and 80NSSC22K1934.
+R.S.
+acknowledges funding from the CAS-ANID grant number CAS220016.
+C.R.
+acknowledges support from Fondecyt Regular grant 1230345, ANID BASAL project FB210003, and the China-Chile joint research fund.